PLEASE HELP!Write the first five terms of the five terms of the sequence defined by an=-2n-5
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

A1=(-2*1)-5
A2=(-2*2)-5
A3=(-2*3)-5
A4=(-2*4)-5
A5=(-2*5)-5

Answer by stanbon(75887)

Write the first five terms of the five terms of the sequence
defined by an=-2n-5
-----------
a(1) = -2*1-5 = -7
a(2) = -2*2-5 = -9
a(3) = -2*3-5 = -6-5 = -11
a(4) = -2*4-5 = -13
a(5) = -2*5-5 = -15
==================
Cheers,
Stan H.